An 12 g ice cube at -19 C is put into a Thermos flask containing 100 cm23 of water at 20 C. By how much has the entropy of the cube-water system changed when a final equilibrium state is reached? The specific heat of ice is 2200 J/kg K and that of liquid water is 4187 3/kg K. The heat of fusion of water is 333 x 103 J/kg. specific heat of ice is 2200 J/kg K and that of liquid water is 4187 J/kg K. The heat of fusion of water is 333 x 103 J/kg Number 0.245 Unity J/KAn 12 g ice cube at -19?C is put into a Thermos flask containing 100 cm3 of water at 20?C. By how much has the entropy of the cube-water system changed when a final equilibrium state is reached? The specific heat of ice is 2200 J/kg K and that of liquid water is 4187 J/kg K. The heat of fusion of water is 333 × 103 J/kg.
Answer #1

let the final temperature be T

by heat conservation

heat given=heat taken

(0.012*2200*19)+(0.012*333000)+(0.012*4187*T)=0.1*4187*(20-T)

solving for T we get

T=8.26 C=281.26 K

heat exchanged=Q=4912.9 J

entropy change=Q/T=17.47 J/K